non-small-caps shapes: upshape: This Is Arbitrary Text ⟹^{upshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{itshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{slshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{scshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{noscshape} This Is Arbitrary Text

itshape: This Is Arbitrary Text ⟹^{upshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{itshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{slshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{scshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{noscshape} This Is Arbitrary Text

slshape: This Is Arbitrary Text ⟹^{upshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{itshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{slshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{scshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{noscshape} This Is Arbitrary Text

small-caps shapes: upshape: This Is Arbitrary Text ⟹^{upshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{itshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{slshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{scshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{noscshape} This Is Arbitrary Text

itshape: This Is Arbitrary Text ⟹^{upshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{itshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{slshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{scshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{noscshape} This Is Arbitrary Text

slshape: This Is Arbitrary Text ⟹^{upshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{itshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{slshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{scshape} This Is Arbitrary Text This Is Arbitrary Text ⟹^{noscshape} This Is Arbitrary Text